OpenStudy (anonymous):

9^x = 3^(1-2x) What is x? Explain to me how to do these ridiculous problems with the exponents being all subtracted.

OpenStudy (anonymous):

trick is to write with the same base, then compare exponents

OpenStudy (anonymous):

you have \[9^x\] and \[9=3^2\] so this is \[(3^2)^x=3^{2x}\]

OpenStudy (anonymous):

so the left hand side is \[3^{2x}\] right hand is \[3^{1-2x}\] so \[3^{2x}=3^{1-2x}\] ,making \[2x=1-2x\] solve for x

OpenStudy (anonymous):

point is not the exponents being subtracted, point is making them equal to see what you get for x. you get \[4x=1\] \[x=\frac{1}{4}\]

OpenStudy (anonymous):

You're the best, thank you! Makes much more sense now!

OpenStudy (anonymous):

thx and yw